# Benchmarking Label Noise in Instance Segmentation: Spatial Noise Matters

### v2

@@ 4.2 Qualitative analysis / Confidence
Interestingly, we show in Figure 7 that despite these challenges, the model demonstrates the ability to distinguish classification losses into two distinct Gaussian distributions (top row). Spatial noise on the other hand, is hard to separate across all training procedure (bottom row), highlighting the difficulties in designing algorithms that effectively address spatial label noise. This is particularly challenging since most current methods rely on the premise that losses from clean and noisy instances are separable. For more illustrations of the training procedure and applying existing method for learning with label noise refer to Appendix D.
The insights from these evaluations suggest that future work should focus on enhancing model resilience to mixed noise conditions, potentially by integrating hybrid noise adaptation strategies. Further investigations could also explore the application of these robust models in real-world scenarios, where label imperfections are commonplace.
@@ 5 Discussion
Our study defines label noise and demonstrates that instance segmentation models experience performance impairments with corrupted labels. To track future progress in this challenging setup, we propose the Noisy label segmentation Benchmark which includes easy-to-use benchmarks for manually and weakly annotation errors. We provide evidence that performance on our benchmarks yields reduced model effectiveness. Our results also show why standard approaches to dealing with label noise do not lead to improved performance. We have provided a script that can apply the benchmarks to any segmentation dataset; this requires no additional labeling or model tuning. Additionally, our experiments indicate that there is still room for improvement, and it remains to be seen whether the most promising LNL techniques will require architectural modifications, data augmentation schemes, modifications to the loss function, or a combination of these strategies.
Further exploration and a deeper understanding of the effects of label noise on long-tailed class distributions, as studied in the LVIS dataset [11], as well as the impact of self-supervised pre-training [47], are necessary. Preliminary investigations of the former are presented in Appendix D.
We encourage researchers to expand the benchmark with more severe levels of noise to evaluate harder cases of label corruption. Since our benchmark is open-source, we welcome new ideas for introducing spatial noises and look forward to contributions via pull requests.
Limitation: A limitation of our work is that our noise models are inspired by observations on real datasets, making it important to acknowledge that we are adding "one more straw" to already noisy datasets. We strongly believe that proper evaluations require relabeling, similar to the approach by Beyer et al. [2], which could yield even more accurate algorithms for dealing with noisy labels.
Appendix A Additional Experiments
@@
To further validate our hypothesis and noise design choices, we conducted additional experiments. As presented in Table 3, we evaluated the traditional symmetric and asymmetric class noise on instance segmentation using MASK-RCNN with two different backbones to assess the resulting performance degradation.
Models/Labels | Clean | Sym 20 | Sym 50 | Sym 80 | Asym 40 |
Mask RCNN R50 FPN | 38/34.6 | 35.5/31.9 | 32.2/29.2 | 22.5/20.2 | 34.6/31.4 |
Mask RCNN R101 FPN | 40.1/36.2 | 37.5/33.6 | 34.5/31 | 25.2/22.7 | 36.8/33.2 |
@@
Foreground-maskrcnn50 | bbox | segm | boundry |
@@
In addition to evaluating the benchmark itself, we extended our analysis to include the impact on object detection performance. Specifically, we examined the scores, as presented in Table 5. This table highlights the detrimental effects of spatial label noise on bounding box quality, in addition to the previously discussed impacts on mask quality. By analyzing the , we aim to demonstrate the broader implications of our noise design choices, showing that spatial noise not only affects segmentation masks but also significantly degrades the performance of object detection tasks. This comprehensive evaluation underscores the robustness of our benchmark in assessing the performance degradation across different aspects of instance segmentation and object detection.
Dataset | Model | Clean | Easy | Mid | Hard |
COCO-N | M-RCNN (R50) | 38 | 35.4 | 34.3 | 33.4 |
M-RCNN (R101) | 40.1 | 37.4 | 36.5 | 35.2 | |
M2F (R50) | 45.7 | 42.2 | 43.7 | 44.7 | |
M2F (Swin-S) | 49.3 | 47.9 | 47.1 | 45.7 | |
YOLACT (R50) | 30.8 | 29.2 | 28.2 | 27.7 | |
Cityscapes-N | M-RCNN (R50) | 41.5 | 35.7 | 32.8 | 31.2 |
M-RCNN (R101) | 39.8 | 32.8 | 29.6 | 26.8 | |
COCO-WAN | M-RCNN (R50) | 36.3 | 34.1 | 25.5 | 22.4 |
@@
Dataset | Clean | Easy | Mid | Hard | ||||
LVIS-N | 22.8 | 22.1 | 15.5 | 14.3 | 17.7 | 13 | 13.3 | 11.2 |
Appendix B Implementation Details
@@ B.2 Datasets
Appendix C Noise levels
The asymmetric noisy class label is defined as:
In our implementation, each spatial noise mask get equal probability, this we can simplify the noised mask equation to define it as:
In Table 8 we showcase what are the mAP and boundary mAP comparing the original annotations and their noisy counterparts. The top line shows the morphological operations that compose the scale spatial annotation noise and the bottom one the chosen noises we widraw from to create our noisy labels benchmark.
We elaborate on the levels of noise in the benchmark in Table 7, presenting the percentage of class symmetric noise, independent mask spatial noise levels, and intensity by kernels used or approximation degree of redundancy of the polygon as described in the paper.
Figure 9 illustrates masks with increasing intensities of each type of spatial noise. We observed similar behaviors in real datasets, where manual annotations indicate Scale and Approximation errors. Localization errors, even at high intensities, correlate more with datasets that collected annotations using a model.
Level Class Noise Spatial Noise Intensity Easy Low (20%) Low (20%) , , Medium Moderate (30%) Moderate (25%) , , Hard High (40%) High (30%) , ,
C.1 Hardware details
MS-COCO based experiments (include both COCO and LVIS) conducted on local machine with 4 Nvidia RTX A6000 or 4 Nvidia RTX 3090, ranging from 20 hours (Mask-RCNN with R50) to 7 days (Mask2Former with SWIN transformer beckbone), training for 12 epochs for all models except YOLACT that trained for 50 epochs. Cityscapes experiments conducted on local machine with one instance of Nvidia RTX 3090, training for 12 epocs for about 12 hours. All experiments use the default configs from MMDetection [5].
Appendix D Learning with Noisy Labels
@@
Model | Clean | Easy | Mid | Hard | ||||
CE | 34.6 | 20.6 | 31.8 | 18.9 | 30.3 | 17.5 | 28.4 | 16.3 |
SCE | 32.5 | 19.5 | 32.1 | 18.9 | 30.8 | 17.8 | 28.8 | 16.4 |
Appendix E Biases of Self-annotating Datasets

### v3

@@ 4.2 Confidence and Loss Analysis
We examine the model’s ability to distinguish noisy from clean annotations. Figure 10 shows two experiments: in the first, 40% of instances contain class noise; in the second, 40% have medium-level spatial noise. Under class noise, the model’s classification losses form two roughly distinct Gaussian distributions, suggesting partial separation of clean and noisy samples. By contrast, when spatial noise is introduced, the losses remain intermixed throughout training. This highlights the challenge of boundary-level label errors for methods relying on loss-based separation. Further experimental details and additional results on learning with noisy labels appear in the supplementary.
These findings point to the need for future strategies that address both class and spatial noise simultaneously. Potential directions include hybrid noise-adaptation techniques and more robust architectures, particularly for real-world applications where label imperfections are unavoidable.
In Figure 8, we compare the mAP and boundary mAP of original vs. noisy annotations. The top row illustrates the morphological operations used for scale-based spatial distortion, while the bottom row shows the specific noise types we apply in our benchmark.
Our experiments indicate that various architectures and backbones exhibit notable sensitivity to label noise, affecting both mask quality and prediction confidence. As shown in Figure 9, higher noise levels correlate with reduced confidence scores, underscoring the vulnerability of model predictions to annotation accuracy. This effect is further illustrated in Figure 12, where increased noise leads to misclassification, causing the model to generate multiple conflicting predictions for a single instance.
@@ 5 Discussion
Our experiments demonstrate that label noise—whether from imprecise human annotations, automated tools, or weak prompts—can substantially degrade the performance of instance segmentation models. We introduced both synthetic and weakly annotated benchmarks that systematically capture real-world noise patterns, ranging from boundary misalignments to class confusion and missing instances. Even moderate levels of noise can erode confidence in model predictions and lead to notable mAP reductions, highlighting the sensitivity of current architectures to spatial inaccuracies.
In particular, our results show that (1) models trained on large datasets like COCO and Cityscapes are far from robust under moderate noise, exhibiting over 10% drops in mask mAP, (2) scale noise—especially erosions—can severely mislead boundary-based metrics, and (3) while promptable foundation models reduce labeling effort, they also introduce new biases and are not fully immune to noisy prompts or ambiguous object boundaries. These outcomes underscore the gap between current label-noise handling strategies—mostly devised for image classification—and the complexities of segmentation tasks, where spatial quality is paramount.
We hope that the public release of our noise-generation toolkit, along with reproducible benchmarks, will encourage further research into mitigating annotation noise and inspire more resilient models for real-world segmentation.
Supplementary Material
6 Ejection Fraction Analysis in the CAMUS Dataset
The CAMUS dataset [23] provides 2D echocardiographic images along with high-quality, expert-annotated labels of the left ventricle (LV). A critical clinical metric in these annotations is the left ventricle’s ejection fraction (EF), defined as:
where EDV is the end-diastolic volume (i.e., the LV volume at its most dilated state) and ESV is the end-systolic volume (the LV volume at maximal contraction). EF offers a succinct quantification of cardiac pump efficiency: a healthy range is typically considered to be above 50%, while borderline or reduced EF can indicate impaired cardiac function.
Clinical Implications and Risks.
Misestimations of the LV boundary—especially at the end-diastolic or end-systolic frames—can propagate into disproportionate errors in volume computations. Even small annotation noise around the boundary may shift the EF from borderline-normal (e.g., 45%) to a clearly abnormal () or misleadingly normal () reading. Such inaccuracies pose a risk for misdiagnosis or delayed therapeutic intervention, since EF underlies critical clinical decisions, including the prescription of certain medications, lifestyle interventions, or further diagnostic procedures.
Noise-Induced Errors.
Figure 13 (to be added) illustrates how a noisy annotation around the LV boundary at end-diastole can lead to an overestimation or underestimation of EDV. When combined with an equally skewed ESV, the net EF deviation can be clinically significant. We examine morphological dilation of the ESV boundary, along with moderate localization noise in both EDV and ESV, using the “low” noise setup described in the main text.
Evaluation Under Noisy Labels.
We trained a simple convolution-based U-Net model, as described in [23], on both clean and noisy CAMUS annotations, and compared the results in Table 7. Evaluation metrics are Dice Index for segmentation overlap of the left ventricle (LV) at end-systolic (ES) and end-diastolic (ED) frames,EF Error as mean absolute error compared to 2D compute of EF values from the labels in percentage points (p.p), as well as HD (Hausdorff Distance) for boundary alignment.
Training | Dice (%) | EF Error | HD (mm) | |
Data | ES | ED | (p.p.) | (ED frame) |
Clean | 86.9 | 91.1 | 2.1 | 6.3 |
Noisy | 82.1 | 87.5 | 4.5 | 11.25 |
As Table 7 indicates, the model trained on noisy labels tends to yield worse Dice overlap and a higher EF error than when trained on clean labels, underscoring the sensitivity of medical diagnostics to annotation precision. Crucially, this discrepancy demonstrates that even modest boundary errors can propagate into clinically important EF ranges, highlighting the urgency of robust noise-handling strategies in echocardiographic segmentation tasks.
7 Additional Experiments
@@
To further validate our noise design choices and their impact, we conducted additional experiments. As presented in Table 9, we evaluated the traditional symmetric and asymmetric class noise on instance segmentation using MASK-RCNN with two different backbones to assess the resulting performance degradation. “Sym ” refers to symmetric class confusion with probability , while “Asym ” denotes mislabeling concentrated in a smaller set of classes [28, 39].
Dataset Model Clean Easy Medium Hard COCO-N M-RCNN (R50) 20.6 18.9 17.5 16.3 M-RCNN (R101) 22.2 20.4 19.0 17.4 M2F (R50) 30.0 28.6 26.7 23.8 M2F (Swin-S) 32.6 30.9 29.3 26.2 YOLACT (R50) 15.7 14.4 13.5 12.4 Cityscapes-N M-RCNN (R50) 33.4 28.4 24.7 22.8 M-RCNN (R101) 34.3 30.7 29.0 25.4 YOLACT (R50) 16.5 16.5 14.5 13.3
Models/Labels Clean Sym 20% Sym 50% Sym 80% Asym 40% M-RCNN (R50) 38/34.6 35.5/31.9 32.2/29.2 22.5/20.2 34.6/31.4 M-RCNN (R101) 40.1/36.2 37.5/33.6 34.5/31 25.2/22.7 36.8/33.2
@@
Foregound noise | bbox | segm | boundry |
@@
In addition to evaluating the benchmark itself, we extended our analysis to include the impact on object detection performance. Specifically, we examined the and scores, as presented in Tables 8 and11 respectively. This tables highlights the detrimental effects of spatial label noise on the boundaries of the masks, as well as bounding box quality, in addition to the previously discussed impacts on mask quality. By analyzing the , we aim to demonstrate the broader implications of our noise design choices, showing that spatial noise not only affects segmentation masks but also significantly degrades the performance of object detection tasks. This comprehensive evaluation underscores the robustness of our benchmark in assessing the performance degradation across different aspects of instance segmentation and object detection.
Dataset Model Clean Easy Medium Hard COCO-N M-RCNN (R50) 38 35.4 34.3 33.4 M-RCNN (R101) 40.1 37.4 36.5 35.2 M2F (R50) 45.7 42.2 43.7 44.7 M2F (Swin-S) 49.3 47.9 47.1 45.7 YOLACT (R50) 30.8 29.2 28.2 27.7 Cityscapes-N M-RCNN (R50) 41.5 35.7 32.8 31.2 M-RCNN (R101) 39.8 32.8 29.6 26.8 COCO-WAN M-RCNN (R50) 36.3 34.1 25.5 22.4
@@
Dataset Clean Easy Medium Hard LVIS-N 22.8 22.1 15.5 14.3 17.7 13 13.3 11.2
8 Additional Noise Visualizations
Figure 14 presents additional samples from our benchmark under different intensities of spatial label noise. Each row highlights a specific set of distortions—such as boundary approximations or morphological operations—applied to one or more instances. As the noise severity increases from left to right, the object contours become visibly degraded, illustrating the range of realistic annotation errors our benchmark can simulate.
9 Implementation Details
@@ 9.2 Datasets
VIPER
VIPER [34] is a synthetic dataset generated from the GTA V game engine. It provides per-pixel annotations for a broad range of 31 categories in photorealistic urban scenes, making it ideal for benchmarking under controlled conditions. Because VIPER annotations are automatically rendered (rather than hand-labeled), they are virtually free from human annotation errors, allowing precise evaluation of how injected label noise affects segmentation performance.
@@
9.3 Hardware details
MS-COCO based experiments (include both COCO and LVIS) and VIPER conducted on local machine with 4 Nvidia RTX A6000 or 4 Nvidia RTX 3090, ranging from 20 hours (Mask-RCNN with R50) to 7 days (Mask2Former with SWIN transformer beckbone), training for 12 epochs for all models except YOLACT that trained for 50 epochs. Cityscapes experiments conducted on local machine with one instance of Nvidia RTX 3090, training for 12 epocs for about 12 hours. All experiments use the default configs from MMDetection [5].
10 Learning with Noisy Labels
@@
Loss Clean Easy Mid Hard CE 34.6 20.6 31.8 18.9 30.3 17.5 28.4 16.3 SCE 32.5 19.5 32.1 18.9 30.8 17.8 28.8 16.4
11 SAM Finetune with label noise
Since for weakly supervision annotations we heavly relay on SAM [21], we exemine how noise in prompt effect the model itself in two setups, zero-shot, that corespond to the quality of the masks produced by sam, and fine-tuning, as a popular paradigm of using SAM for a downstream application. For the zero-shot, we prompt SAM with the grounded bounding boxes of the validation set of COCO as well as noisy boxes with the COCO-N hard type of noise on the validation annotations. For fine-tuned, we exemine fine tuning with both clean and noisy COCO-N hard annotations masks. Table 14, shows both mIoU and F1 scores of the masks produced by SAM, showing that the quality of masks can be increased when fine-tuned, compared with zero shot training with high quality prompts. Fine-tuning with noisy annotations however, is less sever, when prompting with cerfully designed prompts, compared to noisy prompts. Our findings suggest that the quality of prompts are fur more important then the qua
Annotations | Clean | COCO-N Hard | ||
Method | ||||
Zero-shot | 79.78 | 87.49 | 67.99 | 63.30 |
Fine-tune | 79.91 | 78.6 | 77.47 | 76.18 |
12 Biases of Self-annotating Datasets